# Frequency Decomposition-Driven Unsupervised Domain Adaptation for Remote Sensing Image Semantic Segmentation

### v1

@@ 5.5 Results and Discussions
Non-DA (Source-Only)
ResUNet++
42.87
0.10
7.36
3.55
43.27
34.14
19.43
12.11
DANN
63.70
42.50
40.26
58.21
62.10
54.53
53.35
37.00
AdasegNet
ResNet101
73.23
46.23
33.88
59.48
71.61
58.38
56.89
41.27
Advent
ResNet101
71.13
46.62
31.70
66.76
70.95
57.30
57.43
41.90
GANAI
ResNet101
64.97
38.79
45.29
67.45
67.17
57.02
56.73
40.58
TriADA
ResNet101
73.95
38.06
47.87
70.46
72.55
61.34
60.58
44.99
CCAGAN
ResNet101
76.45
44.86
47.28
72.82
71.26
61.68
62.53
46.87
MBATA-GAN
ResNet101
81.54
30.75
58.14
69.83
75.37
65.82
63.13
48.42
GLGAN
Swin-Base
82.77
60.46
65.19
77.27
76.14
71.15
72.37
57.34
Swin-Large
85.65
62.08
69.26
81.37
76.09
73.63
74.89
60.58
FD-GLGAN
Swin-Base
85.65 65.61 71.02
79.43
79.66 75.54 76.27 62.17
Non-DA (Supervised)
ResUNet++
92.08
74.91
80.13
90.79
89.35
84.69
85.46
75.19
Figure 11: Task 3: qualitative results of the adaptation V-IRRG to P-RGB with the size of $2048\times 2048$. (a) RGB images, (b) Ground Truth, (c) Source-only, (d) AdasegNet, (e) Advent, (f) GANAI, (g) TriADA, (h) CCAGAN, (i) MBATA-GAN, (j) GLGAN, (k) FD-GLGAN and (l) Supervised method.
Furthermore, it is interesting to observe that the FD-GLGAN improved the accuracy for Low Vegetation by $20.51\%$, but suffered from performance degradation of $0.32\%$ for Trees. This observation could be explained by the fact the trees in Potsdam are mostly sparse deciduous trees, while those in Vaihingen have denser foliage and canopy, as shown in Fig. 5. As a result, Trees and Low Vegetation in Potsdam share similar spectral characteristics with Low Vegetation in Vaihingen.
Finally, Fig. 10 illustrates some segmentation examples obtained in Task $2$ by various methods. It is clearly observed that the “Source-only” method suffered from the worst performance. Furthermore, despite the fact that existing UDA methods showed improvements of different degrees, they incurred deficiencies in the aspects of providing clear boundaries and differentiating visually similar categories. By leveraging its powerful global-local feature extraction capability and frequency decomposition frequency technology, the proposed FD-GLGAN achieved segmentation performance close to that by the supervised non-DA method.
Task 3 (V-IRRG to P-RGB): Task $3$ suffers from a domain shift problem similar to Task $2$, except that its source domain has a smaller data size than the target domain. Tabel 6 summarizes the results in Task $3$. From Table 6, it is observed that the proposed FD-GLGAN achieved the highest OA of $75.54\%$, mF1 of $76.27\%$, and mIoU of $62.17\%$, which stands for an improvement of $18.24\%$, $18.84\%$, and $20.27\%$ as compared to Advent, respectively. Furthermore, the segmentation accuracy on Tree and Low vegetation in Task $3$ has been substantially improved as compared to Task $2$. This is because these two categories in Vaihingen show much different spectral characteristics. Some segmentation samples are presented in Fig. 11 for visual inspection. In sharp contrast to other UDA methods, the proposed methods provided the most accurate semantic information with the smoothest boundaries.
In summary, the extensive experiments demonstrated that the proposed GLGAN and FD-GLGAN have the capability of overcoming domain shifts, leading to significantly improved semantic segmentation performance on remote sensing images.
5.6 Ablation Studies
Considering that we have proposed GLGAN based on GLTB, and HLFD based on decomposition technology, we conducted three sets of ablation studies to illustrate the necessity of each designation.
1) Global-Local information: To verify the effectiveness of the proposed backbone GLGAN, eight ablation experiments were performed as shown in Table 7. The baseline shown in the first row was obtained with a purely convolutional model composed of an encoder (Resnet101), a decoder (ASPP) and a CNN-based discriminator whereas the second row shows the case for the baseline with Resnet101 being replaced by the Swin Transformer for feature extraction. Furthermore, the third row shows the case in which the ASPP in the baseline was replaced by the GLTB-based decoder to recover the semantic features and predict the segmentation maps, while the CNN-based discriminator in the baseline was substituted with GLDis in the fourth row. The next three rows, i.e. the fifth row to the seventh row, are pairwise combinations of the Swin Transformer-based encoder, the GLTB-based decoder and GLDis. Inspection of the first four rows in Table 7 suggests that each proposed component is of great necessity in the proposed GLGAN.
Furthermore, the ablation experiments shown in the second and fifth rows confirmed the importance of the decoder in learning a transferable output space collocated with the Swin Transformer. Comparison of the results reported in the second and sixth rows, the third and seventh rows, and the fifth and last rows suggest that GLDis was able to improve overall segmentation performance, which provides strong evidence to support our motivation to strengthen the discriminator design under the GAN framework. Clearly, GLDis is a better choice than the conventional CNN-based discriminator. Table 7 confirmed that the proposed GLGAN is more efficient in image feature extraction and semantic information recovery by effectively exploiting intra- and inter-domain features.
Table 7: Ablation studies on the structure of GLGAN (%). Bold values are the best
@@ 5.6 Ablation Studies
Table 8: Ablation studies for high/low-frequency of FD-GLGAN (%). HF denotes high-frequency and LF denotes low-frequency. Bold values are the best
@@
2) High/Low-frequency information: To verify the effectiveness of the two branches in HLFD, four ablation experiments were performed as shown in Table 8. The baseline shown in the first row was the results provided by GLGAN. The second row shows the case that only the high-frequency extractor is employed to align the high-frequency information whereas the third row shows the case that only the low-frequency extractor is employed to align the low-frequency information. The last row presents the results achieved by utilizing the complete HLFD. Inspection of the results suggests that in UDA tasks, the separate alignment of high- and low- frequency information is very important when learning cross-domain transferable features. We can also draw an inference that besides sharing similarities in global features, the source and target domains also exhibit resemblances in local features, such as the shape of tree crowns and vehicle outlines. These local characteristics of ground objects can be analyzed from a frequency perspective, thereby offering valuable insights for unlabeled target domains.
3) Multiscale features: Eight ablation experiments were performed as shown in Table 9 to verify the necessity of aligning multiscale features. The baseline shown in the first row was still the results provided by GLGAN. According to the experimental results we obtained, we present the ablation experiments in the order shown in Table 9. Specifically, we investigated the impact of a single scale, ranging from shallow features (scale 1) to deep features (scale 4), before progressively incorporating additional scales. There are two interesting conclusions confirmed by Table 9. Firstly, the alignment of multiscale features is necessary. Secondly, deep features hold greater significance compared to shallow features that is also widely recognized in other works (Chen et al., 2022; Ma et al., 2023b; Liu et al., 2022).
Table 9: Ablation studies on the multiscale features of FD-GLGAN (%). Bold values are the best
@@
In summary, the proposed GLGAN and FD-GLGAN achieved great segmentation performance as compared to other state-of-the-art UDA methods. Both quantitative results and visual inspection confirmed that GLTB-based backbone GLGAN and the frequency decomposition technology can largely bridge the segmentation performance gap between UDA methods and supervised methods.
Figure 12: Parameter analysis for high/low-frequency of FD-GLGAN.
5.7 Parameter Analysis
@@
The sensitivity of hyper-parameter $\lambda^{f}$ is analyzed, as presented in Fig. 12. This parameter is designed to regulate the contribution derived from frequency decomposition. Inspection of the results in Fig. 12 indicates that a smaller value of $\lambda^{f}$ diminishes the importance of aligning high/low-frequency information, whereas a larger value may exaggerate the significance of the frequency alignment. Consequently, setting $\lambda^{f}\geq 0.05$ results in notable performance degradation. Conversely, performance demonstrates less sensitivity to $\lambda^{f}$ within the range $\lambda^{f}\in\left[0.008,0.02\right]$. Therefore, we set $\lambda^{f}=0.01$ in our experiments.
Figure 13: Visualization of feature maps before and after decomposition. The first row are the input images from the source and the target domains. The second to the fifth row corresponds to the four scales. Source domain feature maps: (a) before decomposition, (b) low-frequency and (c) high-frequency. Target domain feature maps: (d) before decomposition, (e) low-frequency and (f) high-frequency.
5.8 Visualization of Feature Decomposition
@@
In this section, we will investigate the characteristics of the proposed multiscale HLFD modules by visually inspecting the heatmaps before and after frequency decomposition. Specifically, multiscale feature maps are first resized to a uniform image size, followed by averaging across the channel dimension. Subsequently, pseudo-color processing is applied. The results are depicted in Fig. 13. Upon inspecting the feature maps before and after decomposition in both the source and target domains, distinctions become apparent between low- and high-frequency information: low-frequency features possess global characteristics, whereas high-frequency features appear more discrete and localized. In particular, even the fourth scale (high-level) feature maps show significant differences. It fully demonstrates that our HLFD modules successfully decompose high- and low-frequency information.
@@
Comparing feature maps between the source and target domains reveals that high-frequency features cross-domain share great similarities as well as the low-frequency features. It suggests that the model’s capacity to decompose features acquired from the labeled source domain can be successfully applied to the target domain. In other words, the model has learned domain-invariant features from both high- and low- frequency, which correspond to the similarity of spatial details across domains that we emphasized in Sec. 1 and the global context similarity that the existing works mainly explores. Our work has proved that both of these information are necessary for cross-domain alignment in the UDA method.
5.9 Model Complexity Analysis

### v2

@@ IV-E Results and Discussions
Non-DA (Source-Only) ResUNet++ 42.87 0.10 7.36 3.55 43.27 34.14 19.43 12.11
DANN - 63.70 42.50 40.26 58.21 62.10 54.53 53.35 37.00
AdasegNet ResNet101 73.23 46.23 33.88 59.48 71.61 58.38 56.89 41.27
Advent ResNet101 71.13 46.62 31.70 66.76 70.95 57.30 57.43 41.90
GANAI ResNet101 64.97 38.79 45.29 67.45 67.17 57.02 56.73 40.58
TriADA ResNet101 73.95 38.06 47.87 70.46 72.55 61.34 60.58 44.99
CCAGAN ResNet101 76.45 44.86 47.28 72.82 71.26 61.68 62.53 46.87
MBATA-GAN ResNet101 81.54 30.75 58.14 69.83 75.37 65.82 63.13 48.42
MASNet ResNet101 77.89 44.87 61.89 76.04 73.97 66.42 66.93 51.51
DSSFNet ResNet101 79.82 50.62 64.68 76.86 75.47 68.87 69.49 54.22
GLGAN Swin-Base 82.77 60.46 65.19 77.27 76.14 71.15 72.37 57.34
Swin-Large 85.65 62.08 69.26 81.37 76.09 73.63 74.89 60.58
De-GLGAN Swin-Base 85.65 65.61 71.02 79.43 79.66 75.54 76.27 62.17
Non-DA (Supervised) ResUNet++ 92.08 74.91 80.13 90.79 89.35 84.69 85.46 75.19
Task 3 (V-IRRG to P-RGB): Task $3$ suffers from a domain shift problem similar to Task $2$, except that its source domain has a smaller data size than the target domain. From Table VI, it is observed that De-GLGAN achieved the highest OA of $75.54\%$, mF1 of $76.27\%$, and mIoU of $62.17\%$, which stands for an improvement of $18.24\%$, $18.84\%$, and $20.27\%$ as compared to Advent, respectively. Furthermore, the segmentation accuracy on Tree and Low vegetation in Task $3$ has been substantially improved as compared to Task $2$. This is because these two categories in Vaihingen show much different spectral characteristics, providing more transferable features. Some segmentation samples are presented in Fig. 12 for visual inspection. In sharp contrast to other UDA methods, the proposed methods provided the most accurate semantic information with the smoothest boundaries.
Figure 12: Task 3: qualitative results of the adaptation V-IRRG to P-RGB with the size of $1024\times 1024$, where the two samples are distinguished by subscripts. (a) RGB images, (b) Ground Truth, (c) AdasegNet, (d) Advent, (e) TriADA, (f) CCAGAN, (g) MBATA-GAN (h) MASNet, (i) DSSFNet and (j) De-GLGAN. Purple boxes are marked to highlight the differences.
Task 4 (Rural to Urban): Inspection of Table VII suggests that the proposed De-GLGAN achieved the highest OA, mF1 and mIoU among all the unsupervised methods. Compared with the baseline, the proposed De-GLGAN achieved great improvements in all performance metrics. Specifically, the OA, mF1 and mIoU attained were $70.01\%$, $70.93\%$ and $55.50\%$, respectively, which amounts to an improvement of $6.33\%$, $6.32\%$ and $7.4\%$ as compared to Advent. Notably, the proposed methods attained great performance in two categories, including $59.53\%$ for Background and $74.49\%$ for Building. These two categories account for more than half of the distribution of Urban scene as presented in Fig. 9(c). It revealed that our approaches effectively learn the transferable features of the dominant object categories in the rural scene, which can aid in identifying these categories in the urban scene. Fig. 13 shows the semantic segmentation results for this task. Visual inspection of Fig. 13 reveals that our method still has fewer random prediction points, maintaining the integrity of the ground objects.
TABLE VII: Task 4: Quantitative comparison of F1 (%) on the adaptation of Rural to Urban.
Method Backbone Bac. Bui. Road Wat. Bar. For. Agr. OA mF1 mIoU
DANN - 53.38 60.11 56.59 74.63 58.51 64.80 65.19 61.71 61.89 45.14
AdasegNet ResNet101 54.76 62.77 60.84 80.75 62.28 63.23 65.57 63.02 64.31 47.87
Advent ResNet101 56.40 65.53 62.52 77.98 57.72 65.53 66.60 63.68 64.61 48.09
GANAI ResNet101 55.54 61.04 54.76 75.42 65.31 65.01 67.31 62.98 63.48 46.86
TriADA ResNet101 57.05 63.35 65.17 80.56 65.36 67.16 71.53 66.24 67.17 50.97
CCAGAN ResNet101 56.91 67.45 65.68 83.00 65.10 65.93 74.93 67.78 68.43 52.55
MBATA-GAN ResNet101 58.66 64.95 65.90 82.54 64.35 67.72 74.18 67.93 68.47 52.52
MASNet ResNet101 59.49 71.28 64.15 83.36 63.34 69.97 75.99 69.52 69.65 53.98
DSSFNet ResNet101 59.25 73.95 65.84 84.44 64.12 70.74 72.89 68.90 70.18 54.60
GLGAN Swin-Base 58.06 73.11 69.99 84.09 59.70 70.66 73.64 69.13 69.89 54.32
Swin-Large 57.83 74.45 67.90 84.57 63.08 69.09 75.68 69.50 70.37 54.80
De-GLGAN Swin-Base 59.53 74.49 68.80 84.66 63.71 70.50 74.83 70.01 70.93 55.50
TABLE VIII: Task 5: Quantitative comparison of F1 (%) on the adaptation of Urban to Rural.
Method Backbone Bac. Bui. Road Wat. Bar. For. Agr. OA mF1 mIoU
DANN - 67.80 49.19 24.80 57.99 26.04 7.30 28.43 55.04 37.36 24.89
AdasegNet ResNet101 66.47 62.36 34.59 68.49 27.85 15.05 17.04 54.85 41.69 28.82
Advent ResNet101 68.49 62.01 32.89 54.88 33.65 5.44 39.92 57.82 42.47 28.93
GANAI ResNet101 67.75 63.08 30.20 72.08 20.00 7.62 26.21 56.64 40.99 28.80
TriADA ResNet101 69.91 66.06 53.17 57.67 18.76 22.61 54.99 62.58 49.02 34.40
CCAGAN ResNet101 70.34 65.28 45.76 68.60 21.97 11.72 55.33 63.45 48.43 34.48
MBATA-GAN ResNet101 70.00 64.77 49.93 67.46 16.03 7.19 51.53 62.12 46.70 33.27
MASNet ResNet101 68.71 67.13 41.53 55.06 19.11 12.10 49.46 60.18 44.73 30.99
DSSFNet ResNet101 70.88 61.88 45.79 65.88 23.94 29.88 52.54 63.01 50.11 35.01
GLGAN Swin-Base 69.68 63.65 40.80 69.87 25.72 6.65 44.15 60.90 45.79 32.27
Swin-Large 69.74 67.80 51.73 71.91 15.10 8.20 48.63 62.20 47.59 34.34
De-GLGAN Swin-Base 71.48 70.96 49.92 69.12 23.79 33.60 56.83 65.22 53.67 38.58
Figure 13: Task 4: qualitative results of the adaptation Rural to Urban with the size of $1024\times 1024$, where the two samples are distinguished by subscripts. (a) RGB images, (b) Ground Truth, (c) AdasegNet, (d) Advent, (e) TriADA, (f) CCAGAN, (g) MBATA-GAN (h) MASNet, (i) DSSFNet and (j) De-GLGAN. Purple boxes are marked to highlight the differences.
Figure 14: Task 5: qualitative results of the adaptation Urban to Rural with the size of $1024\times 1024$, where the two samples are distinguished by subscripts. (a) RGB images, (b) Ground Truth, (c) AdasegNet, (d) Advent, (e) TriADA, (f) CCAGAN, (g) MBATA-GAN (h) MASNet, (i) DSSFNet and (j) De-GLGAN. Purple boxes are marked to highlight the differences.
Task 5 (Urban to Rural): Compared with Task $4$, Task $5$ is more challenging since the minority categories in the Urban scene, including Water, Forest and Agriculture, are precisely the categories of objects with complex characteristics. Specifically, compared with objects such as Building and Road, they have more complicated and variable boundaries and spatial relationships. From Table VIII, it is observed that the proposed De-GLGAN achieved the highest OA of $65.22\%$, mF1 of $53.67\%$, and mIoU of $38.58\%$, which stands for an improvement of $7.4\%$, $11.2\%$, and $9.65\%$ as compared to Advent, respectively. Furthermore, the segmentation accuracy on Building, Forest and Agriculture in Task $5$ has been greatly improved. Observing the results of De-GLGAN and GLGAN, we can see that the frequency decomposition technique has a powerful feature extraction capability for this difficult category. By comparing the results of GLGAN and De-GLGAN, it is evident that the frequency decomposition technique excels in feature extraction for the challenging categories. The HLFD module can focus on both local details and global semantics, enabling the extraction of transferable features from a minimal number of category samples. The visualized segmentation samples shown in Fig. 14 indicated that De-GLGAN was more accurate in predicting Agriculture and Forest. It further demonstrates that our method can effectively learn transferable features from a limited number of samples.
The comparative results reveal a performance gap between our method and other methods. This is because the proposed framework improves both the generator and discriminator of the GAN framework. In summary, the extensive experiments demonstrated that the proposed GLGAN and De-GLGAN have the capability of overcoming domain shifts, leading to significantly improved semantic segmentation performance on different remote sensing scenarios.
IV-F Ablation Studies
Given that the proposed De-GLGAN incorporates both GLTB and HLFD, we conducted three sets of ablation studies to demonstrate the importance of each component.
1) Global-Local information: To verify the effectiveness of the proposed backbone GLGAN, eight ablation experiments were performed as shown in Table IX. The baseline shown in the first row was obtained with a purely convolutional model composed of an encoder (ResNet101), a decoder (ASPP) and a CNN-based discriminator, whereas the second row shows the case for the baseline with ResNet101 being replaced by the Swin Transformer for feature extraction. Furthermore, the third row shows the case in which the ASPP in the baseline was replaced by the GLTB-based decoder to recover the semantic features and predict the segmentation maps, while the CNN-based discriminator in the baseline was substituted with GLDis in the fourth row. The next three rows, i.e., the fifth row to the seventh row, are pairwise combinations of the Swin Transformer-based encoder, the GLTB-based decoder and GLDis. Inspection of the first four rows in Table IX suggests that each proposed component is of great necessity in the proposed GLGAN. Notably, the result in the second row shows that Swin Transformer only contributes a small part of the superior performance of GLGAN. Furthermore, the ablation experiments shown in the second and fifth rows confirmed the importance of the decoder in learning a transferable output space collocated with the Swin Transformer. Comparison of the results reported in the second and sixth rows, the third and seventh rows, and the fifth and last rows suggest that GLDis was able to improve overall segmentation performance, which provides strong evidence to support our motivation to strengthen the discriminator design under the GAN framework. Therefore, the outstanding performance of GLGAN benefits from the combined effect of all the aforementioned modules. Clearly, GLDis is a better choice than the conventional CNN-based discriminator. Table IX confirmed that the proposed GLGAN is more efficient in image feature extraction and semantic information recovery by effectively exploiting intra- and inter-domain features.
TABLE IX: Ablation studies on the structure of GLGAN (%). Bold values are the best.
@@ IV-F Ablation Studies
TABLE X: Ablation studies for high/low-frequency of De-GLGAN (%). Bold values are the best.
@@
2) High/Low-frequency information: To verify the effectiveness of the two branches in HLFD, four ablation experiments were performed, as shown in Table X. The baseline shown in the first row was the results provided by GLGAN. The second row shows the case in which only the high-frequency extractor is employed to align the high-frequency information, whereas the third row shows the case in which only the low-frequency extractor is employed to align the low-frequency information. The last row presents the results achieved by utilizing the complete HLFD. Inspection of the results suggests that in UDA tasks, the separate alignment of high- and low-frequency information is very important when learning cross-domain transferable features. We can also draw an inference that besides sharing similarities in global features, the source and target domains also exhibit resemblances in local features, such as the shape of tree crowns and vehicle outlines. These local characteristics of ground objects can be analyzed from a frequency perspective, thereby offering valuable insights for unlabeled target domains.
3) Multiscale features: Eight ablation experiments were performed, as shown in Table XI to verify the necessity of aligning multiscale features. The baseline shown in the first row still included the results provided by GLGAN. According to the experimental results we obtained, we present the ablation experiments in the order shown in Table XI. Specifically, we investigated the impact of a single scale, ranging from shallow features (scale 1) to deep features (scale 4), before progressively incorporating additional scales. There are two interesting conclusions confirmed by Table XI. Firstly, the alignment of multiscale features is necessary. Secondly, deep features hold greater significance compared to shallow features, which has been also reported in the literature [32, 21, 51]. In summary, the proposed GLGAN and De-GLGAN achieved great segmentation performance as compared to other state-of-the-art UDA methods. Both quantitative results and visual inspection confirmed that GLTB-based backbone GLGAN and the frequency decomposition technology can largely bridge the segmentation performance gap between UDA methods and supervised methods.
TABLE XI: Ablation studies on the multiscale features of De-GLGAN (%). Bold values are the best.
@@
IV-G Parameter Analysis
@@
In this section, the sensitivity of hyper-parameter $\lambda^{f}$ is analyzed. This parameter is designed to regulate the contribution derived from frequency decomposition. Inspection of the results presented in Fig. 15 indicates that a smaller value of $\lambda^{f}$ diminishes the importance of aligning high/low-frequency information, whereas a larger value may exaggerate the significance of the frequency alignment. Consequently, setting $\lambda^{f}>0.05$ results in notable performance degradation. Conversely, both performance demonstrates less sensitivity to $\lambda^{f}$ within the range $\lambda^{f}\in\left[0.008,0.02\right]$. Therefore, we set $\lambda^{f}=0.01$ in our experiments, which performs effectively across the five UDA tasks, suggesting its strong applicability.
Figure 15: Parameter analysis for high/low-frequency of De-GLGAN on (a) P-IRRG to V-IRRG and (b) Rural to Urban. The result indicates that as $\lambda^{f}$ increases, the impact of feature decomposition gradually improves until excessively large values lead to overemphasis.
IV-H Visualization of Feature Decomposition
@@
In this section, we will investigate the characteristics of the proposed multiscale HLFD modules by visually inspecting the heatmaps before and after the frequency decomposition. Specifically, multiscale feature maps are first resized to a uniform image size, followed by averaging across all channels. Subsequently, pseudo-color processing is applied. The results are depicted in Fig. 16. Upon inspecting the feature maps before and after the decomposition in both the source and target domains, distinctions become apparent between low- and high-frequency information: low-frequency features possess global characteristics, whereas high-frequency features appear more discrete and localized. In particular, even the fourth scale (high-level) feature maps show significant differences. It fully demonstrates that our HLFD modules successfully decompose high- and low-frequency information.
Figure 16: Visualization of feature maps before and after decomposition. The first row displays the input images from the source and the target domains. The second to the fifth row corresponds to the four scales. Source domain feature maps: (a) before decomposition, (b) low-frequency and (c) high-frequency. Target domain feature maps: (d) before decomposition, (e) low-frequency and (f) high-frequency.
@@
Comparison on the feature maps between the source and target domains reveals that cross-domain high-frequency features share great similarities as well as the cross-domain low-frequency features. It suggests that the model’s capacity to decompose features acquired from the labeled source domain can be successfully applied to the target domain. In other words, the model has learned domain-invariant features from both high- and low-frequency, which correspond to the similarity of spatial details across domains that we emphasized in Sec. I and the global context similarity that the existing works explored. Our work has proved that both of these components are necessary for cross-domain alignment in the UDA method.
IV-I Model Complexity Analysis